4. How much time is required for 5mg of 22Na (T


1


2 = 2.60 yrs) to reduce to 1 mg?

1
Expert's answer
2021-12-27T08:19:15-0500

"N=N_0\times(\frac{1}{2})^{\frac{t}{t_{1/2}}}"


"1\ mg=5\ mg\times(\frac{1}{2})^{\frac{t}{2.60\ years}}"


Solving for t, we get t = 6.04 years.


Answer: 6.04 years